z= xy+f(x^2+y^2+z^2)



1
Expert's answer
2021-05-04T12:32:59-0400

"\displaystyle
z = xy + f(x^2 + y^2 + z^2)\\

\frac{\partial z}{\partial x} = y + 2xf(x^2 + y^2 + z^2)\\

\frac{\partial z}{\partial y} = x + 2yf(x^2 + y^2 + z^2)\\


y\frac{\partial z}{\partial x} = y^2 + 2xyf(x^2 + y^2 + z^2)\\

x\frac{\partial z}{\partial y} = x^2 + 2xyf(x^2 + y^2 + z^2)\\


y\frac{\partial z}{\partial x} - x\frac{\partial z}{\partial y} = y^2 - x^2"